Answers Provided: Without excel-how do I calculate them with a calculator?-again answers provided

Month Orders
January 130
February 138
March 152
April 160
May 173
June 184
July 191
August 196
September 201
October 211
November 222
December 237

What is the 4-period moving average for May? =145.00

What is the 4-period moving average for June? =155.75

What is the 4-period moving average for July? =167.25

What is the 4-period moving average for August?=177.00

What is the 4-period moving average for September?=186.00

What is the 4-period moving average for October?=193.00

What is the 4-period moving average for November?199.75

What is the 4-period moving average for December?=207.50

What is the 4-period moving average for next January?=217.75
Answer #1

Let actual orders for period t be denoted by At

Let the forecast for period t be denoted by Ft

For a 4 period moving average, Ft = (At-1 + At-2 + At-3 + At-4)/4

Calculating using the above formula -

Month (t) Orders (A) Forecast (F)
January 130
February 138
March 152
April 160
May 173 =(130+138+152+160)/4=145
June 184 =(138+152+160+173)/4=155.75
July 191 =(152+160+173+184)/4=167.25
August 196 =(160+173+184+191)/4=177
September 201 =(173+184+191+196)/4=186
October 211 =(184+191+196+201)/4=193
November 222 =(191+196+201+211)/4=199.75
December 237 =(196+201+211+222)/4=207.5
January =(201+211+222+237)/4=217.75